what is the smallest of three positive consecutive odd integers if the product of the second and third integers is 63?
Found 2 solutions by mananth, algebrahouse.com:
Answer by mananth(16946)

let the consequtive odd integers be n , n+2, n+4
...
(n+2)(n+4)=63
n^2+6n+8=63
n^2+6n-55=0
n^2+11n-5n-55=0
n(n+11)-5(n+11)=0
(n+11)(n-5)=0
n=5 positive integer
...
the numbers are 5,7,9

Answer by algebrahouse.com(1659)

"what is the smallest of three positive consecutive odd integers if the product of the second and third integers is 63?"

x = first odd integer
x + 2 = second odd integer
x + 4 = 3rd odd integer

(x + 2)(x + 4) = 63 {product of second and third is 63}
x^2 + 6x + 8 = 63 {used foil method}
x^2 + 6x - 55 = 0 {subtracted 63 from both sides}
(x + 11)(x - 5) = 0 {factored into two binomials}
x + 11 = 0 or x - 5 = 0 {set each factor equal to 0}
x = 5 {first positive odd integer}
x + 2 = 7 {second positive odd integer}
x + 4 = 9 {third positive odd integer}

5 would be the smallest positive odd integer
www.algebrahouse.com